### v5

# On the Hidden Mystery of OCR in Large Multimodal Models

Abstract
Large models have recently played a dominant role in natural language processing and multimodal vision-language learning. However, their effectiveness in text-related visual tasks remains relatively unexplored. In this paper, we conducted a comprehensive evaluation of Large Multimodal Models, such as GPT4V and Gemini, in various text-related visual tasks including Text Recognition, Scene Text-Centric Visual Question Answering (VQA), Document-Oriented VQA, Key Information Extraction (KIE), and Handwritten Mathematical Expression Recognition (HMER). To facilitate the assessment of Optical Character Recognition (OCR) capabilities in Large Multimodal Models, we propose OCRBench, a comprehensive evaluation benchmark. Our study encompasses 29 datasets, making it the most comprehensive OCR evaluation benchmark available. Furthermore, our study reveals both the strengths and weaknesses of these models, particularly in handling multilingual text, handwritten text, non-semantic text, and mathematical expression recognition. Most importantly, the baseline results showcased in this study could provide a foundational framework for the conception and assessment of innovative strategies targeted at enhancing zero-shot multimodal techniques. The evaluation pipeline and benchmark are available at https://github.com/Yuliang-Liu/MultimodalOCR.
Keywords Large Multimodal Model, OCR, Text Recognition, Scene Text-Centric VQA, Document-Oriented VQA, Key Information Extraction, Handwritten Mathematical Expression Recognition
1 Introduction
The advent of large models has unlocked a wealth of potentials in the realm of advanced computing. In recent years, there has been an explosion in the development of large language models (LLM) such as ChatGPT [1] and GPT-4 [2], giving rise to extraordinary applications in zero-shot task transfer to many new real-world scenarios. The success of proprietary LLMs has stimulated tremendous interest in developing open-source LLMs. Among them, LLaMA [3] is an open-source LLM that matches the performance of GPT-3, followed by Alpaca [4], Vicuna [5], GPT-4-LLM [6] to improve the LLM’s alignment ability to follow human instruction, reporting impressive performance compared with proprietary LLMs.
The success of large models has also been extended to the multimodal vision-language space [7], leading to a line of research on large multimodal models (LMM), including contrastive learning [8, 9, 10, 11] and generative modeling [2, 12, 13, 14, 15]. Surprisingly, Liu et al. [15] show that LMM exhibits excellent zero-shot OCR performance in the wild, without explicitly training on the OCR domain-specific data. Understanding the efficacy of LMM in handling text-related visual tasks is pivotal, given their potential to infer context from multiple data sources, such as text and images. Despite this advantage, these models may face challenges when dealing with complex relationships between different data types due to their general training on web-scale data. Recognizing these limitations could guide improvements in multimodal methodologies and inspire the creation of more robust models that can handle text-related tasks more efficiently. Additionally, this knowledge can open up novel applications in areas such as digital marketing or social media analysis, where understanding the interplay between textual and visual content in the images is crucial.
To this end, we undertake a comprehensive study on 14 LMMs by evaluating their OCR ability on five representative tasks: text recognition, Scene Text-Centric VQA, Document-Oriented VQA, key information extraction and handwritten mathematical expression recognition. Our findings indicate that even state-of-the-art large multimodal models, such as  [16] and GPT4V [17], encounter difficulties in recognizing handwritten mathematical expressions. LMMs continue to struggle with blurry text images, handwritten text, multilingual text. Moreover, they heavily rely on semantic understanding to recognize words, often favoring common words over random letter sequences. Consequently, these findings highlight even the most powerful LMM leaves a gap to match the domain-specific methods in various traditional text tasks, and even worse for more complicated tasks. However, it highlights the vast opportunity to improve the OCR ability of LMM by adapting them to the OCR domains.
2 Related Work
2.1 Large Multimodal Models
The remarkable success of Large Language Models (LLMs) has paved the way for the development of large multimodal models, which combine pretrained visual models with LLMs to enable their visual capabilities. BLIP2 [18] introduces the Querying Transformer (Q-Former) as a means to bridge the gap between vision and language models. Flamingo [13] and OpenFlamingo [19] enhance a frozen pretrained LLM by incorporating novel gated cross-attention-dense layers, enabling conditioning on visual inputs. LLaVA [15] pioneers the use of GPT-4 to generate multimodal instruction-following data. Other works, such as [20, 21], also focus on aligning the vision module and LLM for improved multimodal understanding. Additionally, [22, 23] emphasize modality collaboration for image and text. LLaVAR [24] collects training data with rich text and uses a higher-resolution CLIP as the visual encoder to enhance LLaVA’s OCR ability. BLIVA [25] combines instruction-aware and global visual features to capture richer image information. MiniGPT4V2 [26] uses unique identifiers for different tasks when training the model to better distinguish each task instruction effortlessly. UniDoc [27] performs unified multimodal instruction tuning on large-scale instruction following datasets and leverages the beneficial interactions among tasks to enhance the performance of each individual task. Docpedia [28] directly processes visual input in the frequency domain rather than the pixel space. Monkey [29] enhanced the LMM’s ability to perceive details at a low cost by the generated detailed caption and a high-resolution model architecture.
2.2 Benchmarks for Large Multimodal Models
With the ongoing advancements in Large Multimodal Models, the question of how to effectively evaluate their performance has emerged. [15, 30, 22] introduces GPT-4 or human evaluation to assess the output of LMMs. MMBench [31] evalute LMM using multiple-choice questions across various dimensions of ability. MME [32] measures both perception and cognition abilities on True or False question. However, their testing on OCR data is limited. Additionally, the True/False question or multiple-choice question cannot accurately assess the OCR ability to recognize words in an image. In this paper, we conducted an extensive study on various LMMs for five prominent text-related tasks. To facilitate the evaluation of LMMs’ OCR ability, we also present OCRBench, a collection of 1000 manually filtered and corrected question-answer pairs on five representative text-related tasks.
Figure 1: Visualization results of the five tasks. ‘[Q]’ represents question, ‘[G]’ represents ground truth, ‘[P]’ represents prediction generated by LMM.
3 Experiments
3.1 Evaluation Metric and Evaluation Dataset
The responses generated by LMM often include many explanatory terms, so the exact matching approach or Average Normalized Levenshtein Similarity(ANLS) [33] used in the original dataset are not suitable for evaluating LMM in zero-shot scenarios. We have defined a unified and simple evaluation criterion for all datasets, which is to determine whether the ground truth (GT) is present in the output of the LMM. To reduce false positives, we filter out questions that have answers containing fewer than 4 symbols from all datasets. Additionally, we choose 3000 question instances from some large datasets for testing purposes.
Text Recognition: We evaluate LMM using widely-adopted OCR text recognition datasets, including (1) Regular Text Recognition: IIIT5K [34], SVT [35], IC13 [36]; (2)Irregular Text Recognition: IC15 [37], SVTP [38], CT80 [39], COCOText(COCO) [40], SCUT-CTW1500 (CTW) [41], Total-Text (TT) [42]; (3) Occlusion Scene Text [43], encompassing weakly occluded scene text (WOST) and heavily occluded scene text (HOST); (4) Artistic Text Recognition: WordArt [44]; (5) Handwirtten Text Recognition: IAM [45]; (6) Chinese Text Recognition: ReCTS [46]; (7) Handwritten Digit String Recognition: ORAND-CAR-2014(CAR-A) [47]; (8) Non-Semantic Text(NST) and Semantic Text(ST): LMMs primarily rely on semantic understanding to recognize words. In the experiments, we find that the LMMs have poor recognition performance on character combinations that lacked semantics. To confirm this, we create two datasets: Semantic Text (ST) and Non-Semantic Text (NST) using the IIIT5k dictionary. The ST dataset consists of 3000 images with words from the IIIT5K dictionary, while the NST dataset contains the same words but with shuffled characters without semantics. For english words, we employ the consistent prompt "what is written in the image?". For Chinese words in ReCTS dataset, we adapt the prompt to "What are the Chinese characters in the image?". For handwritten digit string, we utilize the prompt "what is the number in the image?".
Scene Text-Centric VQA: We test LMMs on STVQA [48], TextVQA [49], OCRVQA [50] and ESTVQA [51]. Scene Text Visual Question Answering (STVQA) consists of 31,000+ questions across 23,000+ images collected from various public datasets. TextVQA dataset comprises 45,000+ questions on 28,000+ images sampled from specific OpenImages dataset categories expected to contain text. OCRVQA features over 1 million question-answer pairs spanning 207,000+ book cover images. ESTVQA contains 20757 images along with 15056 English questions and 13006 Chinese questions. We have divided the ESTVQA dataset into ESTVQA(CN) and ESTVQA(EN), which specifically include questions and answers in Chinese or English respectively.
Document-Oriented VQA: We assess these LMMs on DocVQA[52], InfographicVQA [53] and ChartQA [54]. DocVQA is a large-scale dataset with 12,767 document images of diverse types and content, and over 50,000 questions and answers. InfographicVQA is a diverse collection of infographics that includes 5,485 images and a total of 30,035 questions. ChartQA includes a total of 9,608 human-written questions covering 4,804 charts, as well as 23,111 questions generated from human-written chart summaries on 17,141 charts.
Key Information Extraction: We conduct experiments on SROIE [55], FUNSD [56] and POIE [57]. SROIE contains 1000 complete scanned receipt images for OCR and key information extraction competitions. In this competition, it is required to extract the company, date, address, and total expenditure information based on the receipts. FUNSD dataset consists of 199 real, fully annotated, scanned forms that may contain noise. POIE consists of camera images from Nutrition Facts label of products in English and 3,000 images with 111,155 text instances are collected. KIE dataset requires to extract key-value pairs in the image. To enable LMMs to accurately extract the correct value for a given key in the KIE dataset, we employ manual prompt design. For SROIE dataset, we utilize the following prompts to assist LMMs in generating the respective values for "company", "date", "address", and "total": "what is the name of the company that issued this receipt?", "when was this receipt issued?", "where was this receipt issued?", and "what is the total amount of this receipt?". Additionally, to retrieve the corresponding value for a given key in FUNSD and POIE, we utilize the prompt "What is the value for ‘{key}’?"
Handwritten Mathematical Expression Recognition(HMER): We evaluate on HME100K [58], which consists of 74,502 images for training and 24,607 images for testing, with 245 symbol classes. During evaluation, we use the prompt "Please write out the expression of the formula in the image using LaTeX format.".
Method Regular Irregular Occluded Artistic Handwritten Chinese Digits Semantic
IIIT5K SVT IC13 IC15 SVTP CT80 COCO CTW TT HOST WOST WordArt IAM ReCTS ORAND NST ST
BLIP2-OPT-6.7B 79.1 86.7 83.4 71.2 78.8 76.5 51.4 62.4 68.7 59.8 69.4 68.4 32.9 0 1.2 13.0 82.6
mPLUG-Owl 81.1 84.3 85.9 67.5 73.9 81.4 52.3 69.2 74.4 49.7 62.7 72.1 34.8 0 13.5 44.7 92.3
InstructBLIP 86.3 92.0 86.8 80.9 85.6 86.3 62.6 70.8 77.9 67.8 78.8 73.7 42.6 0 18.4 25.3 89.7
LLaVAR 84.0 87.6 87.7 79.4 84.0 84.5 61.9 69.5 75.6 61.1 71.9 67.1 49.4 0 9.8 36.2 86.5
BLIVA 86.5 90.6 87.3 80.9 87.7 86.7 64.8 71.2 78.1 67.7 77.6 73.7 45.1 0 13.8 20.4 89.4
mPLUG-Owl2 80.9 69.6 79.8 53.9 53.5 74.8 52 59.1 60.9 32.5 50.6 60.6 23.8 0 9.9 48.2 93.9
LLaVA1.5-7B 84.2 85.7 86.4 71.9 79.8 82.7 55.6 66.8 73.2 61.4 70.6 68.7 55.4 0 10.4 15.2 85.3
UniDoc 91.9 89.2 90.9 78 80.3 88.2 64.1 75.3 78.2 52.4 68.5 - - - - - -
Monkey 83.7 75.1 85.4 53.4 58.4 73.9 43.5 64.5 64.6 43.3 54.9 67.7 30.3 13.1 29.1 56.5 95.5
Supervised-SOTA 96.6 93.0 96.7 85.7 89.3 89.9 64.4 78.6 80.1 73.1 81.6 72.5 91.2 94.8 95.5 95.4 100.0
Table 1: Text recognition results. Bold black digits indicate the best result, while blue signifies the second best.
Method Scene Text-Centric VQA Document-Oriented VQA. KIE HMER
STVQA TextVQA OCRVQA ESTVQA(EN) ESTVQA(CN) DocVQA InfoVQA ChartQA FUNSD SROIE POIE HME100K
BLIP2-OPT-6.7B 20.9 23.5 9.7 40.7 0 3.2 11.3 3.4 0.2 0.1 0.3 0
mPLUG-Owl 30.5 34 21.1 52.7 0 7.4 20 7.9 0.5 1.7 2.5 0.1
InstructBLIP 27.4 29.1 41.3 48.6 0.1 4.5 16.4 5.3 0.2 0.6 1 0
LLaVAR 39.2 41.8 24 58.2 0 12.3 16.5 12.2 0.5 5.2 5.9 0
BLIVA 32.1 33.3 50.7 51.2 0.2 5.8 23.6 8.7 0.2 0.7 2.1 0.1
mPLUG-Owl2 49.8 53.9 58.7 68.6 4.9 17.9 18.9 19.4 1.4 3.2 9.9 0
LLaVA1.5-7B 38.1 38.7 58.1 52.3 0 8.5 14.7 9.3 0.2 1.7 2.5 0
UniDoc 35.2 46.2 36.8 - - 7.7 14.7 10.9 1 2.9 5.1 0.4
DocPedia 45.5 60.2 57.2 - - 47.1 15.2 46.9 29.9 21.4 39.9 -
Monkey 54.7 64.3 64.4 71 42.6 50.1 25.8 54.0 24.1 41.9 19.9 0.2
Supervised-SOTA 69.6 73.7 68.1 43.3 43.3 90.16 36.8 70.5 93.1 98.7 79.5 64.3
Table 2: Results of Scene Text-Centric VQA, Document-Oriented VQA KIE and HMER. Bold black digits indicate the best result, while blue signifies the second best. Since the Supervised-SOTA on ESTVQA dataset does not provide separate results on Chinese and English data, the average performance 42.3 is used as a reference.
3.2 Results
The results of text recognition are shown in Tab. 1. LMMs achieve comparable performance to state-of-the-art supervised models in recognizing regular text, irregular text, occluded text, and artistic text. Even in the WordArt dataset, which primarily features challenging artistic text, InstructBLIP2 and BLIVA surpass the performance of supervised state-of-the-art model. However, LMMs perform poorly in recognizing handwritten text, Chinese text, handwritten strings, and non-semantic text. In tasks such as Scene Text-Centric VQA, Document-Oriented VQA, and KIE, LMMs with smaller input resolutions consistently yield poorer results compared to those with larger input sizes. This is because texts often contain intricate structures and can vary in size, requiring the LMMs to capture fine details. The results are shown in Tab. 2 LMMs struggle to complete the complex task of recognizing handwritten mathematical expressions. Through extensively analyzing the results, we qualitatively summarize the limitations of LMMs in text tasks:
  • Semantic-Reliance. LMMs primarily rely on semantic understanding to recognize words. In our experiments, we observed that LMMs exhibit poor recognition performance when it comes to character combinations that lack semantic meaning. Specifically, when we altered the order of letters in each word, the accuracy of LMMs on the NST dataset decreased by an average of 57.0% compared to the ST dataset, while the SOTA method for scene text recognition only drops by around 4.6%. We believe this is because the SOTA method for scene text recognition directly recognizes each character, and semantic information is just used to assist the recognition process, while LMMs primarily rely on semantic understanding to recognize words. This finding is further supported by the low accuracy observed on the ORAND dataset. As shown in Fig. 1, LMM successfully identified the word "Message," but incorrectly recognized "egaesMs," which is a rearranged version of the word "Message."
  • Handwritten Text. LMMs may encounter difficulties in accurately recognizing handwritten text due to various reasons, including the resemblances in shapes between handwritten letters and numbers. Handwritten text often appears incomplete or blurry due to factors like fast writing speed, irregular handwriting, or low-quality paper. On average, LMMs perform 51.9% worse than supervised state-of-the-art model in this task.
  • Multilingual Text. The notable performance gap observed in LMMs on ReCTS, ESTVQA(En), and ESTVQA(Ch) highlights their limited proficiency in the Chinese language, as indicated in Tab. 1 and Tab. 2. Among the LMMs, accurately recognizing Chinese words or responding to Chinese questions proves to be a challenge. Monkey outperforms other LMMs in Chinese scenarios due to its LLM and visual encoder being trained on a substantial amount of Chinese data.
  • Fine-grain Perception. Currently, most LMMs have their input images limited to a resolution of 224 x 224, consistent with the input size of the visual encoder used in their architecture. However, high-resolution input images can capture more details from the image. Due to the limited input resolution of LMMs such as BLIP2, their capacity to extract fine-grained information in tasks like Scene Text-Centric VQA, Document-Oriented VQA, and KIE is constrained. However, LMMs such as Monkey, which can support a resolution of 1344×896, exhibit improved performance in these specific tasks.
  • HMER. LMMs struggle to recognize handwritten mathematical expressions due to the presence of messy handwritten characters, complex spatial structures, and the indirect LaTeX representation. The lack of training data for the handwritten mathematical expression recognition task is also a crucial factor.
Method Recog. $VQA^{S}$ $VQA^{D}$ KIE HMER Final Score
Gemini 215 174 128 134 8 659
GPT4V 167 163 146 160 9 645
Monkey 174 161 91 88 0 514
mPLUG-Owl2 153 153 41 19 0 366
LLaVAR 186 122 25 13 0 346
LLaVA1.5-13B 176 129 19 7 0 331
LLaVA1.5-7B 160 117 15 5 0 297
mPLUG-Owl 172 104 18 3 0 297
BLIVA 165 103 22 1 0 291
InstructBLIP 168 93 14 1 0 276
BLIP2-OPT-6.7B 154 71 10 0 0 235
MiniGPT4V2 124 29 4 0 0 157
Table 3: Results of LMMs on OCRBench. Recog. represents text recognition, $VQA^{S}$ represents Scene Text-Centric VQA, $VQA^{D}$ represents Document-Oriented VQA. Bold black digits indicate the best result.
3.3 OCRBench
Evaluating multiple datasets can be time-consuming, and the inaccurate annotations in some datasets make accuracy-based evaluation less precise. Realizing these limitations, we have developed OCRBench to facilitate the accurate and convenient evaluation of LMMs’ OCR capabilities. OCRBench consists of five components: text recognition, Scene Text-Centric VQA, Document-Oriented VQA, KIE, and HMER. It includes 1000 question-answer pairs, and for the KIE task, we added the prompt "Answer this question using the text in the image directly." to restrict the model’s response format. The specific composition of OCRBench is shown in appendix. To ensure a more accurate evaluation, we manually verified and corrected incorrect answers for the 1000 question-answer pairs, providing alternative correct answer candidates. The final results are presented in the Tab. 3, with Gemini achieving the highest score and GPT4V securing the second position. It is important to note that due to rigorous safety reviews by OPEN AI, GPT4V refused to provide results for 84 images in OCRBench. Monkey demonstrated OCR capabilities that were only inferior to GPT4V and Gemini. From the test results, we can observe that even state-of-the-art models like GPT4V and Gemini struggle with the HMER task. Furthermore, they also face challenges in handling unclear images, handwritten text, Non-Semantic text and adhering to task instructions. As shown in Fig. 2 (g), even when explicitly requested to answer using the text found in the image, Gemini consistently interprets "02/02/2018" as "2 February 2018".
Figure 2: Some erroneous results from Gemini and GPT-4V. ‘[Q]’ represents question, ‘[G]’ represents ground truth. (a) showcases results on blurry image, (b) presents results on handwritten text, (c) and (d) showcase results on Non-Semantic text, (e) and (f) showcase results on HMER, (g) displays instances where the task instructions are not adhered to.
4 Discussions
Why LMMs work for OCR? While we offer some analysis based on the results, the question as to why these multimodal models can deliver acceptable performance on OCR tasks remains challenging to conclusively explain. One plausible explanation for the success of multimodal models in Optical Character Recognition (OCR) tasks lies in the training of multimodal models (similar to CLIP), which we believe includes some OCR data. However, unsupervised text-image pairs in the training data cannot compete with fully supervised data.
From a network perspective, the pre-trained models of the visual encoder and Large Language Model (LLM) already demonstrate a strong grasp of their respective domain data, each operating within their structured feature space. These LMMs bridge visual and language information using components such as a linear projection layer, that functions as a step of visual tokenization. This process aligns visual tokens within the pre-trained language model’s word embedding space, resulting in visual embeddings that closely mirror their corresponding word embeddings. This alignment facilitates text recognition, and the LLM subsequently presents this OCR data to users in a generative manner.
Future research could benefit from an ablation analysis to understand the impact of the volume of multimodal training data on OCR performance.
Future Directions: Despite demonstrating promise, current LMMs struggle with complex tasks and cannot rival domain-specific methods in traditional text tasks. To improve, these models should focus on refining their capabilities in detecting fine-grained features in images, enhancing their understanding of individual character shapes, and building multilingual data for training.
Recent work such as the ICL-D3IE [59], introduces an in-context learning framework leveraging large language models like GPT-3 and ChatGPT for document information extraction. This approach effectively navigates the modality and task gaps in this field, extracting challenging and distinct segments from difficult training documents, designing relationship demonstrations for enhanced positional understanding, and incorporating formatting demonstrations for answer extraction. Despite its notable success across various datasets and settings, ICL-D3IE encounters challenges in the in-domain setting on CORD [60]. This result points to the potential of large language models in managing tasks involving visually intricate documents and encourages the development of novel, minimally supervised OCR techniques.
Moreover, a study by Li et al. [61], evaluates ChatGPT across seven detailed information extraction tasks. This evaluation reveals that ChatGPT performs well in OpenIE settings, generates high-quality responses, shows a tendency towards overconfidence, and maintains strong fidelity to original texts. However, this study focused exclusively on pure text, without venturing into visually rich document texts within the OCR domain. Future research should therefore explore ChatGPT’s performance in these visually rich contexts.
The application of LMMs in specialized domains is also a crucial field, and recent developments have shown their vast potential. One such model, developed by Google, is Med-PaLM2 [62], which is fine-tuned on medical knowledge from PaLM2. It is the first language model to perform an expert level on US Medical Licensing Examination (USMLE) style questions and can analyze patients’ conditions through medical images, including plain films and mammograms. Google claims that it has approached the performance of clinician experts. Additionally, LLaVA-Med [63] attempts to extend multimodal instruction-tuning to the biomedical domain and demonstrates excellent chat abilities with domain knowledge. These studies highlight the potential of LMMs in vertical domain tasks. Further exploration of their applications in other domains, such as gaming and education, is also warranted.
Ultimately, these developments and future research directions could potentially pave the way for multimodal models that can more efficiently handle complex tasks like OCR, expanding the application range of LMM.
5 Conclusion
This paper has presented an extensive study on performance of LMM on OCR tasks, including text recognition, Scene Text-Centric, Document-Oriented VQA, KIE and HMER. Our quantitative assessment reveals that LMM can achieve promising results, especially in text recognition, even attaining SOTA performance in some datasets. However, significant gaps persist compared to domain-specific supervised methods, suggesting that specialized techniques tailored to each task are still essential, as the latter uses much less computational resources and data. However, the emergence of LMM does present promising directions for advancements in OCR techniques. It stimulates an intuitive query for potential improvements: after training on a larger amount of text-related data, can LMM surpass supervised methods? We hope this work provides readers with valuable insights into the application of LMM for future exploration in OCR technology development. Importantly, the baseline results presented in this paper can serve as a foundation for the development and evaluation of novel ideas aimed at improving text-related multimodal techniques.
6 Acknowledgements
Thanks for the help from Hongliang Li, Wenwen Yu, Yang Liu, Mingxin Huang, Dezhi Peng, Mingyu Liu and Mingrui Chen.
Appendix A Summary of the Evaluation Benchmarks.
Tab. 4 presents the detailed composition of OCRBench. The Text Recognition task consists of a total of 300 images, including Regular Text Recognition, Irregular Text Recognition, Artistic Text Recognition, Handwriting Recognition, Digit String Recognition, and Non-Semantic Text Recognition, with 50 images each. The Scene Text-centric VQA task includes 200 questions, with 50 questions each from STVQA, TextVQA, ESTVQA(EN), and OCRVQA datasets. The Doc-oriented VQA task comprises 200 questions, with 50 questions each from DocVQA, ChartQA(Aug.), ChartQA(Hum.), and InfoVQA datasets. The Key Information Extraction task consists of 200 questions, with 67 questions from SROIE, 66 questions from FUNSD, and 67 questions from POIE. The Handwritten mathematical expression recognition task includes 100 images from the HME100k dataset. OCRBench encompasses a total of 1000 questions across these five tasks, and all answers have been manually filtered and corrected.
Task Dataset Smaples
Text Recognition Regular Text Recognition 50
Irregular Text Recognition 50
Artistic Text Recognition 50
Handwriting Recognition 50
Digit String Recognition 50
Non-Semantic Text Recognition 50
Scene Text-centric VQA STVQA 50
TextVQA 50
ESTVQA(EN) 50
OCRVQA 50
Doc-oriented VQA DocVQA 50
ChartQA(Aug.) 50
ChartQA(Hum.) 50
InfoVQA 50
Key Information Extraction SROIE 67
FUNSD 66
POIE 67
HMER HME100k 100
All - 1000
Table 4: Details of the OCRBench.
Appendix B Summary of the Models.
Tab. 5 displays the details of the testing models, with most LMMs having their resolution limited to an input size of 224 or 336 due to the visual module. As a result, they perform poorly on Doc-oriented VQA and KIE tasks. Conversely, models with higher resolutions exhibit better performance on these tasks.
Models Language Model Input Resolution
BLIP2-OPT-6.7B OPT-6.7B [64] 224
mPLUG-Owl LLaMA-2 7B [65] 224
InstructBLIP Vicuna-7B [5] 224
LLaVAR Vicuna-7B [5] 336
BLIVA Vicuna-7B [5] 224
mPLUG-Owl2 LLaMA-2 7B [65] 448
LLaVA1.5-7B Vicuna-7B [5] 336
LLaVA1.5-13B Vicuna-13B [5] 336
MiniGPT4V2 Llama2-Chat-7B [65] 224
Monkey Qwen-7B [66] 896
Unidoc Vicuna [5] 336
DocPedia Vicuna [5] 2560
GPT4V gpt-4-vision-preview
Gemini gemini-pro-vision
Table 5: Details of the models.
Appendix C Supervised SOTA
State-of-the-art (SOTA) in widely-adopted OCR text recognition datasets is achieved by PARSeq [67], which utilizes Permutation Language modeling for enhanced contextual information and unifies context-free non-AR and context-aware AR inference. We train Parseq on the ST [68] and MJ [69] synthetic datasets, and test it directly on real datasets. The ReCTS dataset is not commonly used, therefore, we chose TPS-ResNet, which has the highest ranking on the ICDAR ranking table along with a published article, as the SOTA. The model employs Thin-plate-spline (TPS) based Spatial Transformer Network (STN) to normalize the input text images, followed by a ResNet-based feature extractor and BiLSTM for text recognition. Although the evaluation metric for ReCTS is Normalized Edit Distance, we still use it as a reference. AttentionHTR [70] proposes an attention-based sequence-to-sequence model to achieve SOTA on IAM. The word error rate of this method on IAM is 8.76, so we use 91.24 as the corresponding word accuracy. Yu [71] achieves SOTA on ORAND-CAR-2014 by proposing an efficient text line recognition method based on prototype learning with feature-level sliding windows for classification.
For Scene Text Visual Question Answering (STVQA) and Optical Character Recognition Visual Question Answering (OCR-VQA), SOTA is reached by GIT [14]. GIT trains a Generative Image-to-text Transformer which contains one image encoder and one text decoder under a single language modeling task. For TextVQA, Mia [72] achieves SOTA by using T5 [73] for TextVQA task. To align object feature and scene text, Mia is pretrained by masked language modeling(MLM) and relative position prediction(RPP) task. For DocVQA Dataset, BAIDU-DI assembles ERNIE-Layout [74] and DocPrompt (a few-shot model using multi-stage training based on ERNIE-Layout) to achieve SOTA. ESTVQA dataset has not been explored well, Fang [75] achieves SOTA by using a cross-modal attention network to guide the expression of visual and textual features simultaneously while using a transformer decoder to output the results. However, since Fang [75] does not provide separate results on Chinese and English data, the average performance is used as a reference. For InfographicVQA, SOTA is reached by DUBLIN[76]. For ChartQA, Liu[77] achieves SOTA by designing a modality conversion module DEPLOT.
The SOTA on the FUNSD dataset is achieved by ERNIE-Layout [74], which enhances layout knowledge by correcting the reading order in pretraining phase. For SROIE, StrucTexT [78] achieves the SOTA. StrucTexT introduces a segment-token aligned encoder for transformer and is pretrained by Masked Visual Language Modeling task and the new Sentence Length Prediction and Paired Boxes Direction tasks. For POIE, Kuang [57] achieves SOTA by adopting contrastive learning to effectively establish the connections between the tasks of OCR and information extraction. For HME100K, Li [79] proposes a Counting-Aware Network which jointly optimizes HMER and symbol counting tasks to achieve SOTA.